OpenStudy (anonymous):

what two numbers that have the sum of 10 and a difference of 6.

OpenStudy (anonymous):

8+2 = 10, 8-2 = 6

OpenStudy (anonymous):

Let x represent the first number, let y represent the second number x+y = 10 (1) x-y = 6 (2) (1+2) 2x = 16 x = 8 8 + y = 10 y = 10-8 y = 2

OpenStudy (anonymous):

That is how you would officially do it in a difficult question or on a test

OpenStudy (anonymous):

^is that really what you look like?

OpenStudy (anonymous):

Yes, I look like Christian Bale

OpenStudy (anonymous):

BS

OpenStudy (anonymous):

Do you really look like an incline bench?

OpenStudy (anonymous):

Not like I really care who he is...yes I do my arms are giant

OpenStudy (anonymous):

Hahahaha ok bud dyel

OpenStudy (anonymous):

dyel?